OpenStudy (dugalde):

use the half-angle formula to evaluate tan(17pi/12)

OpenStudy (jdoe0001):

\(\bf \cfrac{17}{12}\cdot 2\implies \cfrac{17}{6}\implies \cfrac{12+5}{6}\implies \cfrac{12}{6}+\cfrac{5}{6}\implies 2+\cfrac{5}{6}\qquad thus \\ \quad \\ \cfrac{17\pi }{12}\cdot 2\implies \cfrac{17\pi }{6}\implies \cfrac{12\pi +5\pi }{6}\implies \cfrac{12\pi }{6}+\cfrac{5\pi }{6}\implies 2\pi +\cfrac{5\pi }{6}\qquad thus \\ \quad \\ \cfrac{\frac{17\pi }{12}\cdot 2}{2}\implies \cfrac{17\pi }{12}\qquad thus\qquad \cfrac{2\pi +\frac{5\pi }{6}}{2}\implies \cfrac{17\pi }{12}\)

OpenStudy (jdoe0001):

anyhow... to use the half-angle identity simply multiply the angle given, by 2 and use that in the half-angle identity

OpenStudy (jdoe0001):

\(\bf 2\pi +\cfrac{5\pi }{6}\) means one "go around" and then it lands on \(\bf \cfrac{5\pi }{6}\) so \(\bf 2\pi +\cfrac{5\pi }{6}\) is coterminal with \(\bf \cfrac{5\pi }{6}\)

OpenStudy (dugalde):

so 2pi+(5pi/6)/2 ----17pi/12 is the answer

OpenStudy (jdoe0001):